Savio works for a
cellular phone company. He is paid
$90/day plus $2.00 for each cellular
phone that he sells. Aimee also works
for the cell phone company, but she
makes $120 per day and no extra money
for selling cellular phones.
a) Write an equation to represent the
amount that Savio earns in one day.
Graph the equation.
b) Write an equation to represent the
amount that Aimee earns in one day.
Graph this equation on the same grid
you used in part a).
c) How many cellular phones must
Savio sell in order to make as much
in a day as Aimee?

Answer by Theo(13342)

formula for savio:

y = 2x + 90

formula for aimee:

y = 120

graph of daily compensation for savio and aimee.

$$$

red is compensation for savio.
blue is compensation for aimee.

x represents the number of cell phones sold.
y represents the total compensation.

they break even when savio sells 15 cell phones.

aimee makes 120 no matter how many cell phones she sells.

when savio sells 15 cell phones, he makes 90 + 2 * 15 = 90 + 30 = 120.

if savio sells more than 15 cell phones in a day, he makes more than aimee for that day.